Chuck, a single taxpayer, earns $45,000 in taxable income and $12,700 in interest from an investment in City of Heflin bonds. (Use the U.S tax rate schedule..) (Do not round intermediate calculations. Round "Federal tax" to 2 decimal places.)

Required:

a) How much federal tax will he owe?

b) What is his average tax rate?

c) What is his effective tax rate?

d)What is his current marginal tax rate?

tax rate schedule

2019 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $510,300 $46,628.50 plus 35% of the excess over $204,100
$510,300 $153,798.50 plus 37% of the excess over $510,300

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$           0 $ 19,400 10% of taxable income
$ 19,400 $ 78,950 $1,940 plus 12% of the excess over $19,400
$ 78,950 $168,400 $9,086 plus 22% of the excess over $78,950
$168,400 $321,450 $28,765 plus 24% of the excess over $168,400
$321,450 $408,200 $65,497 plus 32% of the excess over $321,450
$408,200 $612,350 $93,257 plus 35% of the excess over $408,200
$612,350 $164,709.50 plus 37% of the excess over $612,350

Schedule Z-Head of Household

If taxable income is over: But not over: The tax is:
$           0 $ 13,850 10% of taxable income
$ 13,850 $ 52,850 $1,385 plus 12% of the excess over $13,850
$ 52,850 $ 84,200 $6,065 plus 22% of the excess over $52,850
$ 84,200 $160,700 $12,962 plus 24% of the excess over $84,200
$160,700 $204,100 $31,322 plus 32% of the excess over $160,700
$204,100 $510,300 $45,210 plus 35% of the excess over $204,100
$510,300 $152,380 plus 37% of the excess over $510,300

Schedule Y-2-Married Filing Separately

If taxable income is over: But not over: The tax is:
$           0 $    9,700 10% of taxable income
$    9,700 $ 39,475 $970 plus 12% of the excess over $9,700
$ 39,475 $ 84,200 $4,543 plus 22% of the excess over $39,475
$ 84,200 $160,725 $14,382.50 plus 24% of the excess over $84,200
$160,725 $204,100 $32,748.50 plus 32% of the excess over $160,725
$204,100 $306,175 $46,628.50 plus 35% of the excess over $204,100
$306,175 $82,354.75 plus 37% of the excess over $306,175

2) Given the following tax structure

Taxpayer Salary Total tax

mae $12,500 $850

pedro $51,500 ???

A) what is the minimum tax that pedro should pay to make the tax structure vertically equitable with respect to the amount of tax paid ?

minimum tax = ???

b) This would result in what type of tax rate structure Regressive, progressive or proportional ?
Answer #1

Solution:-

Note:- As per HOMEWORKLIB RULES if more than one question is posted than we liable to answer only first question.

a) How much federal tax will he owe:-

Federal Tax $ 5,758.50

Explanation:-

$4,543 + 22% of (45,000 - 39,475)

= 5,758.50

b) What is his average tax rate:-

Average Tax Rate
Total Tax / Taxable Income =
5,758.50 / 45,000 = 12.80 %

c) What is his effective tax rate:-

Effective Tax Rate
Total Tax / Total Income =
5,758.50 / 57,700 = 9.98 %

d) What is his current marginal tax rate:-

Marginal Tax Rate 22 %

Note:- As per HOMEWORKLIB RULES if more than one question is posted than we liable to answer only first question.

Please Rate.